OpenStudy (anonymous):

Solve. Write the answer in simplest form. (Must show work!) 1. 3/4m = 5/6

OpenStudy (mathlegend):

Do you have any ideas @connor_ ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

flip both sides of the equation then multiply both by 1.3 then divide both sides by 4. This is the way 4m/1.3 = 6/5 1.3*(4m/1.3) = (6/5)*1.3 4m =(6/5)*1.3 4m/4 = ((6/5)*1.3)/4 this is your answer put it in a calculator